If f(x)=(x-1)/(x+1), evaluate f^-1((1/c)+1).
hopefully you'll understand that, thank you for your time

Answer by stanbon(75887)

If f(x)=(x-1)/(x+1), evaluate f^-1((1/c)+1
-----
1st: Find f^-1(x):::
y = (x-1)/(x+1)
Interchange x and y to get:
x = (y-1)/(y+1)
Solve for y::
y-1 = x(y+1)
y-xy = x+1
y(1-x) = x+1
---
f^-1(x) = (x+1)/(1-x)
-------
NOw, find f^-1[(1/c)+1)]
----
f^-1[(1/c)+1)] = [(1/c)+2][-1/c]
-----
= (-1/c^2) -(2/c)
-----------------
Cheers,
Stan H.
==============